OpenStudy (anonymous):

!!!!! If f(x) = −2x^5−3x^4−6x^3 / x4, find f′(x)

OpenStudy (anonymous):

I keep getting -10x^4 - 12X^3 - 18x^2 / 4x^3 ....but it's wrong

OpenStudy (anonymous):

I also tried simplifying that, but it didnt work

OpenStudy (anonymous):

What rule do the ratios require?

OpenStudy (anonymous):

i'm not sure what you mean

OpenStudy (anonymous):

I guess i'm missing a big component

OpenStudy (anonymous):

Does quotient rule ring a bell?

OpenStudy (anonymous):

You could also try simplifying the original function then take the derivative or bring it up and use product rule, there's plenty of things you can do with this.

OpenStudy (anonymous):

I simplified and brought it up to -5x^4-4x^3-9x^2

OpenStudy (anonymous):

I do remember the quotient rule!

OpenStudy (anonymous):

but I dont get why the simplified answer didnt work

OpenStudy (anonymous):

that's not the answer.

OpenStudy (campbell_st):

oops should read \[f(x) = -2x - 3 - \frac{18}{4} x^{-2}\]

OpenStudy (campbell_st):

that is the simplified function... now find the 1st derivative

OpenStudy (anonymous):

Ah that looks much better, the derivative should be very easy from here :p

OpenStudy (anonymous):

It just requires the power rule nothing to fancy

OpenStudy (anonymous):

would I multiply 18/4 by -2? sorry, i'm very new to this

OpenStudy (anonymous):

\[\frac{ d }{ dx }x^n = nx^{n-1}\]

OpenStudy (anonymous):

mr. batman!!

OpenStudy (anonymous):

When I use that power rule, the first value (-2x) becomes (-2)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I don't know how to tackle the last value of (18/4) x^ -2 ..... I want to multiple 18/4 by 2 and be left with x^(-2-1 = 3)

OpenStudy (anonymous):

-2 - 3 -18/4 x^-2 -5 - ( -2 * 18/4) x^-3

OpenStudy (anonymous):

(-5) - (-9)x^-3

OpenStudy (anonymous):

(-5) + 9x^-3

OpenStudy (anonymous):

I don't know where I went wrong with the power rule

OpenStudy (campbell_st):

the 1st term is -3x so \[f'(x) = -3 +\frac{9}{x^3}\]

OpenStudy (anonymous):

that's not the answer though...or are you just showing the next step? i'm going to try using the quotient rule...so confused

OpenStudy (campbell_st):

what answer do you have..?

OpenStudy (anonymous):

they don't tell you the answer online, until you get it correct...that's the only way of knowing or waiting until the deadline is over....ill try the quotient rule and see if that works

OpenStudy (anonymous):

you get unlimited chances to guess correct, luckily

OpenStudy (campbell_st):

well here it is \[f(x) = -2x - 3 - \frac{18}{4} x^{-2}\] then \[f'(x) = -2 + \frac{9}{x^3}\]

OpenStudy (anonymous):

I typed that in .. and it was wrong

OpenStudy (anonymous):

hmmm maybe there was a mistake in simplifying

OpenStudy (anonymous):

f(x)= (−2x^5)−(3x^4) − (6x^3) / x4 find f′(x).

OpenStudy (anonymous):

@campbell_st @iambatman

myininaya (myininaya):

hey is that all over x^4 or is just the last term over x^4?

myininaya (myininaya):

\[f(x)=-2x^5-3x^4-\frac{6x^3}{x^4} \text{ or } f(x)=\frac{-2x^5-3x^4-6x^3}{x^4}\]

myininaya (myininaya):

and yes it seems something went wrong with the simplification above

myininaya (myininaya):

\[f(x)=-2x^{5-4}-3x^{4-4}-6x^{3-4}\] I think it was just the last term that might have been written incorrectly

myininaya (myininaya):

try to simplify that yourself and then find the derivative

myininaya (myininaya):

if you want i can check your work

OpenStudy (anonymous):

yes please!!!! ive been on this question for the last 1 hour!!! im going to check what you did right now

OpenStudy (anonymous):

it's all over x^4

OpenStudy (anonymous):

I used the quotient rule and got -2x^8 + 6x^6 / x^8 ...and it was wrong!

myininaya (myininaya):

well do you understand my rewriting of f(x)?

myininaya (myininaya):

\[f(x)=-2x^{5-4}-3x^{4-4}-6x^{3-4} \]

OpenStudy (anonymous):

not really, I dont understand that 4-4 or the 3-4

myininaya (myininaya):

we don't need quotient rule just power rule and constant rule

myininaya (myininaya):

i did top exponent - bottom exponent

OpenStudy (anonymous):

yeah ...can I do it like this:

OpenStudy (anonymous):

-10x^4 - 12x^3 -18x^2 / 4x^3

myininaya (myininaya):

what did you do to get that?

OpenStudy (anonymous):

I brought the 5 down and multiplied it by the -2...

OpenStudy (anonymous):

then took one away from the x and got -10x^4

myininaya (myininaya):

ok i don't know what you did and it doesn't look good and not trying to be mean when i say that but I think you want to look at this the easiest way possible

myininaya (myininaya):

\[f(x)=\frac{-2x^5-3x^4-6x^3}{x^4}=\frac{-2x^5}{x^4}-\frac{3x^4}{x^4}-\frac{6x^3}{x^4}\] wrote as 3 separate fractions

OpenStudy (anonymous):

I was trying to use the quotient rule...but if there's an easier way, that's better

OpenStudy (anonymous):

oh i see where you're talking about

myininaya (myininaya):

\[f(x)=-2x-3-\frac{6}{x}\] do you understand this

OpenStudy (anonymous):

x^0 is one

OpenStudy (anonymous):

that's why it's 3, right?

myininaya (myininaya):

\[f(x)=-2x-3-6x^{-1} \]

myininaya (myininaya):

yes

myininaya (myininaya):

try to differentiate that

myininaya (myininaya):

show me what you get

OpenStudy (anonymous):

ok let me see

myininaya (myininaya):

i will do quotient rule for you after this and show you is lot uglier

OpenStudy (anonymous):

6/x^2 - 2

myininaya (myininaya):

guess what?! :)

myininaya (myininaya):

it looks good

OpenStudy (anonymous):

it took me long to get to that!...how would you differentiate it in an easier way?

OpenStudy (anonymous):

the answer worked!!!!!!!!

myininaya (myininaya):

no that is the easy way

OpenStudy (anonymous):

you're awesome!

myininaya (myininaya):

\[f'(x)=\frac{(-2x^5-3x^4-6x^3)'(x^4)-(x^4)'(-2x^5-3x^4-6x^3)}{(x^4)^2}\] this is the quotient rule setup

OpenStudy (anonymous):

yes, that's exactly what I tried out!

myininaya (myininaya):

one sec my cat is an obstacle i have to work around to type of keyboard be patient with me sorry

OpenStudy (anonymous):

it's okay!!! im patient...I wont go anywhere till I know how to do quotient rule for this question...i tried so many times

myininaya (myininaya):

\[f'(x)=\frac{(-10x^4-12x^3-18x^2)x^4-4x^3(-2x^5-3x^4-6x^3)}{x^8} \\ \frac{-10x^8-12x^7-18x^6+8x^8+12x^7+24x^6}{x^8} \\ \frac{-2x^8+0x^7+6x^6}{x^8} \\ \frac{-2x^8}{x^8}+\frac{0x^7}{x^8}+\frac{6x^6}{x^8} \\ -2+0+\frac{6}{x^2}\]

myininaya (myininaya):

tons uglier but still works

myininaya (myininaya):

you still have to separate the fractions to simplify

OpenStudy (anonymous):

I dont get how you got 0x^7

myininaya (myininaya):

-12x^7+12x^7

myininaya (myininaya):

-12+12=0

OpenStudy (anonymous):

-12x^7 + 12x^7 should just cancel out

myininaya (myininaya):

yes

OpenStudy (anonymous):

thankyou so much!!!! I guess it was just the last step i messed up on!

OpenStudy (anonymous):

needed to simplify the bottom exponent as well!!!!

myininaya (myininaya):

(-12+12)x^7 0x^7 0 <---or yeah this means you can just cancel them out

OpenStudy (anonymous):

thank-you so much!!! where are you from?!! and are you a TA for a university?

myininaya (myininaya):

i retired from teaching a few months ago i'm from southern US

OpenStudy (anonymous):

okay, that makes sense. you explain everything very well

myininaya (myininaya):

I'm horrible at typing Sometimes my sentences come out sorta bad

OpenStudy (anonymous):

nope, I understood everything! you were so patient with me too! I really appreciate it

myininaya (myininaya):

np Good luck on the rest of your work @sharmit I must log off for tonight

OpenStudy (anonymous):

thankyou so much! sweet dreams!